# Yuewei Lin1, Jonathan Koomey4, Xi Yu1, Zhiuha Dong1

### v1

@@ Abstract
Index Terms:
keywords AI training, Sustainable computing, GPU power measurements
I Introduction
The modern era of deep learning required both algorithmic innovation and dedicated hardware. As researchers identified scaling relationships in artificial intelligence (AI) applications between model scale, total training, and performance [kaplan_scaling_2020], computational hardware optimized for these workloads enabled the training of enormous models with tens-to-hundreds of billions of parameters, using thousands of graphics processing units (GPUs). Beginning in late 2022, these steadily improving capabilities captured consumer interest. Cloud infrastructure providers responded with dramatic investment in hardware [rattner_ai_2024], both to meet this demand and in anticipation of further growth.
@@ I Introduction
This work is structured as follows. In Sec. II, we begin with background on the growth of AI infrastructure and the need for accurate power measurements to inform energy use models. Sec. III details our empirical methodology, including the hardware configuration (Sec. IV) tested, the AI training workloads used, and the power measurement approach. We then present and discuss the empirical power demand recorded across three training runs, varying in architecture and memory batching, in Sec. V and VI. Finally, we conclude with a discussion of the limitations of this work, the implications of these findings for data center capacity planning and energy estimation, and directions for future research.
Significant results. The most significant results of our work are summarized as follows:
@@
II Background and Motivations
The commercialization of AI and the substantial hardware investment it requires have drawn increased attention to the energy footprint of these workloads. Coinciding with heightened scrutiny from scholars and the public, companies have reduced the amount of data they release on hardware and metered energy [alben_computing_2024]. This shift is partially due to the maturation of AI applications. As AI transitioned from an academic curiosity to a commercialized product, hardware operation details became a potential source of competitive advantage. Researchers internal to firms conducting training already have access to relevant aggregate data, e.g., their power bill. Several estimates of training energy intensity rely on this proprietary internal power draw data [patel_characterizing_2024, patterson_carbon_2022, wu_sustainable_2022]. Academic researchers instead have to rely on what data they have available, with one common approach to multiply the manufacturer rated power of AI-optimized GPUs or servers [luccioni_estimating_2023] by the reported training time (typically in GPU-hours) of a given model [de_vries_growing_2023, lacoste_quantifying_2019, faiz_llmcarbon_2023]. While this provides valuable bounds, it does not enable granular estimates of node-level power draw, assuming either full utilization of all non-GPU components or neglecting them entirely.
GPU component-level utilization is well specified, in part because most AI-optimized hardware reports it through data center infrastructure management (DCIM) software. In contrast, node-level energy use is often approximated [bouza_how_2023]. There is research grounded in empirical metering of all components or entire nodes [hodak_towards_2019, wang_energy_2023], but these are not generally performed on current generation hardware. There is also some ongoing work characterizing the distinct load profiles of physical modeling and AI workloads [govind_comparing_2023, shankar_trends_2022, zhao_power_2023], though they typically focus on the chip/processor rather than node. Calibrating these estimates with node-level measured power serves to shed further light on the energy use of AI. While AI training and performance benchmarks have been proposed and implemented in the industry [farrell_mlperf_2021], they have not kept pace with commercial system scale-up [alben_computing_2024]. Robust empirical evidence should reflect recent hardware improvements and infrastructure investments. Additionally, there is an opportunity to compare traditional simulation workloads against AI training, as the two are performed on increasingly overlapping hardware configurations.
III Methodology
@@
IV Experimental Setup and Equipment
IV-A Computational Hardware
IV-A1 Cooling and Power Infrastructure
@@ IV-A2 CPU and GPU information
IV-B Training workload and Software Configuration
@@
Image Classification: For the image classification experiments, we used the ResNet-152 [he2016deep] architecture. The CIFAR-10 [Krizhevsky2009LearningML] dataset was employed for these tasks. It consists of 60k $32\times 32$ color images spread across 10 categories (airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck), with 50k images for training and 10k for testing.
@@
VQA: In the VQA experiments, we utilized a Multi-Modal Large Language Model LLaVA [liu2024improved] that combines a visual encoder, CLIP-L-14 [radford2021learning], with a language model decoder, Llama2-13b [touvron2023llama]. The visual encoder processes images, while the Llama2 model, with $\sim$13 billion parameters, focuses on understanding and generating answers based on text. The VQA tasks were trained using a combined dataset of 665k pairs, including image-caption and image-question-answer pairs. These pairs were sourced from COCO [lin2014microsoft], GQA [hudson2019gqa], OCR-VQA [mishra2019ocr], TextVQA [singh2019towards], and Visual Genome [krishna2017visual], each providing rich visual-textual alignment. The COCO [lin2014microsoft] dataset comprises approximately 118k training images, 5k validation images, and 40k test images, with resolutions ranging from $480\times 640$ to $1024\times 768$ pixels, and is primarily used for object detection, segmentation, and image captioning tasks. GQA [hudson2019gqa] includes around 72k images paired with 943k questions, focusing on visual question answering with image sizes typically around $480\times 640$ pixels. OCR-VQA [mishra2019ocr] contains 50k images and questions, designed for text recognition within images. TextVQA [singh2019towards] offers 28,408 images and 45,336 questions, requiring models to interpret text embedded in visual content. Visual Genome [krishna2017visual] provides 108k images, annotated with objects, attributes, and relationships, and is used for tasks like object detection, scene graph generation, and visual reasoning, with image sizes generally between $480\times 640$ and $1024\times 768$ pixels.
@@
Implementation Details: All tasks are implemented using Distributed Data Parallel (DDP) in PyTorch 2.2. For the image classification task, we train the ResNet model from scratch with a batch size of 512 and 4096 for 200 epochs, optimizing the entire model using Stochastic Gradient Descent (SGD) with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of $5\times 10^{-4}$. We employ a cosine annealing learning rate scheduler during training, which adjusts the learning rate based on a cosine function to enhance training performance and convergence. For the Visual Question Answering (VQA) task, we fine-tune the LLaVA [liu2024improved] model on a combined dataset. The LLaVA [liu2024improved] model comprises a visual encoder, a projector, and a large language model (LLM). Specifically, we use CLIP-L-14 [radford2021learning] as the visual encoder, a two-layer MLP with GeLU activation as the projector, and Llama2-13b [touvron2023llama] as the LLM. During training, we keep the visual encoder and projector fixed and train only the LLM. The batch size is set to 128 with learning rate of $2\times 10^{-5}$ and warm up ratio is 0.03 with cosine learning rate scheduler.
@@
GPU Burn: To evaluate the external validity and real-world applicability of our experimental results, we utilized the open-source GPU Burn tool [timonen_wiliccgpu-burn_2024]. A GPU burn is designed to stress test multi-GPU setups by maximizing GPU memory utilization and compute load while validating calculation correctness. The tool forks a process for each GPU, allocates a high percentage of available memory, and performs continuous matrix multiplications using the CUBLAS library. We obtained the gpu-burn source code from the official GitHub repository at github.com/wilicc/gpu-burn and compiled it on our target Linux system following the provided instructions.
@@
IV-C Power Measurement
The GPU power draw was measured using the Weights & Biases (wandb) [wandb] developer interface. The DCIM, provided by NLYTE software, is used to manage, monitor and optimize the data center’s power and cooling infrastructure. Power and current data from the rack mounted PDU is pulled into the DCIM system over the network using communication protocols like SNMP and Modbus. The DCIM collects and processes this data, allowing operators to monitor, analyze, and optimize power usage at the rack level. The DCIM reports real-time data, including metrics like power draw, and alerts for potential issues related to critical infrastructure in the data center.
@@ V Results
TABLE I: Power was sampled at 1-minute intervals over the duration of training. Batch size is measured in images and tokens for the ResNet and Llama models, respectively.
Training Workload Model Task Median Power (kW) Batch Size Duration (min) Total Energy (kWh) Average GPU Utilization (%)
@@
GPU Burn (no CPU) Stress test 7.97 N/A 60 8 100
GPU Burn ($+$CPU) Stress test 8.43 N/A 77 10.5 100
VI Discussion

### v2

# Empirical Measurements of AI Training Power Demand on a GPU-Accelerated Node

@@ Abstract
Keywords: AI training, Sustainable computing, GPU power measurements
1 Introduction
The modern era of deep learning required both algorithmic innovation and dedicated hardware. As researchers identified scaling relationships in artificial intelligence (AI) applications between model scale, total training, and performance [1], computational hardware optimized for these workloads enabled the training of enormous models with tens-to-hundreds of billions of parameters, using thousands of graphics processing units (GPUs). Beginning in late 2022, these steadily improving capabilities captured consumer interest. Cloud infrastructure providers responded with dramatic investment in hardware [2], both to meet this demand and in anticipation of further growth.
@@ 1 Introduction
This work is structured as follows. In Sec. 2, we begin with background on the growth of AI infrastructure and the need for accurate power measurements to inform energy use models. Sec. 3 details our empirical methodology, including the hardware configuration (Sec. 4) tested, the AI training workloads used, and the power measurement approach. We then present and discuss the empirical power demand recorded across three training runs, varying in architecture and memory batching, in Sec. 5 and 6. Finally, we conclude with a discussion of the limitations of this work, the implications of these findings for data center capacity planning and energy estimation, and directions for future research.
1.1 Significant results
The most significant results of our work are summarized as follows:
@@ 1.1 Significant results
2 Background and Motivations
The commercialization of AI and the substantial hardware investment it requires have drawn increased attention to the energy footprint of these workloads. Coinciding with heightened scrutiny from scholars and the public, companies have reduced the amount of data they release on hardware and metered energy [3]. This shift is partially due to the maturation of AI applications. As AI transitioned from an academic curiosity to a commercialized product, hardware operation details became a potential source of competitive advantage. Researchers internal to firms conducting training already have access to relevant aggregate data, e.g., their power bill. Several estimates of training energy intensity rely on this proprietary internal power draw data [4, 5, 6]. Academic researchers instead have to rely on what data they have available, with one common approach to multiply the manufacturer rated power of AI-optimized GPUs or servers [7] by the reported training time (typically in GPU-hours) of a given model [8, 9, 10]. While this provides valuable bounds, it does not enable granular estimates of node-level power draw, assuming either full utilization of all non-GPU components or neglecting them entirely.
GPU component-level utilization is well specified, in part because most AI-optimized hardware reports it through data center infrastructure management (DCIM) software. In contrast, node-level energy use is often approximated [11]. There is research grounded in empirical metering of all components or entire nodes [12, 13], but these are not generally performed on current generation hardware. There is also some ongoing work characterizing the distinct load profiles of physical modeling and AI workloads [14, 15, 16], though they typically focus on the chip/processor rather than node. Calibrating these estimates with node-level measured power serves to shed further light on the energy use of AI. While AI training and performance benchmarks have been proposed and implemented in the industry [17], they have not kept pace with commercial system scale-up [3]. Robust empirical evidence should reflect recent hardware improvements and infrastructure investments. Additionally, there is an opportunity to compare traditional simulation workloads against AI training, as the two are performed on increasingly overlapping hardware configurations.
3 Methodology
@@
4 Experimental Setup and Equipment
4.1 Computational Hardware
4.1.1 Cooling and Power Infrastructure
@@ 4.1.2 CPU and GPU Information
4.2 Training workload and Software Configuration
@@ 4.2.1 Image Classification
For the image classification experiments, we used the ResNet-152 [18] architecture. The CIFAR-10 [19] dataset was employed for these tasks. It consists of 60k $32\times 32$ color images spread across 10 categories (airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck), with 50k images for training and 10k for testing.
4.2.2 VQA
@@
In the VQA experiments, we utilized a Multi-Modal Large Language Model LLaVA [20] that combines a visual encoder, CLIP-L-14 [21], with a language model decoder, Llama2-13b [22]. The visual encoder processes images, while the Llama2 model, with $\sim$13 billion parameters, focuses on understanding and generating answers based on text. The VQA tasks were trained using a combined dataset of 665k pairs, including image-caption and image-question-answer pairs. These pairs were sourced from COCO [23], GQA [24], OCR-VQA [25], TextVQA [26], and Visual Genome [27], each providing rich visual-textual alignment. The COCO [23] dataset comprises approximately 118k training images, 5k validation images, and 40k test images, with resolutions ranging from $480\times 640$ to $1024\times 768$ pixels, and is primarily used for object detection, segmentation, and image captioning tasks. GQA [24] includes around 72k images paired with 943k questions, focusing on visual question answering with image sizes typically around $480\times 640$ pixels. OCR-VQA [25] contains 50k images and questions, designed for text recognition within images. TextVQA [26] offers 28,408 images and 45,336 questions, requiring models to interpret text embedded in visual content. Visual Genome [27] provides 108k images, annotated with objects, attributes, and relationships, and is used for tasks like object detection, scene graph generation, and visual reasoning, with image sizes generally between $480\times 640$ and $1024\times 768$ pixels.
4.2.3 Implementation Details
@@
All tasks are implemented using Distributed Data Parallel (DDP) in PyTorch 2.2. For the image classification task, we train the ResNet model from scratch with a batch size of 512 and 4096 for 200 epochs, optimizing the entire model using Stochastic Gradient Descent (SGD) with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of $5\times 10^{-4}$. We employ a cosine annealing learning rate scheduler during training, which adjusts the learning rate based on a cosine function to enhance training performance and convergence. For the Visual Question Answering (VQA) task, we fine-tune the LLaVA [20] model on a combined dataset. The LLaVA [20] model comprises a visual encoder, a projector, and a large language model (LLM). Specifically, we use CLIP-L-14 [21] as the visual encoder, a two-layer MLP with GeLU activation as the projector, and Llama2-13b [22] as the LLM. During training, we keep the visual encoder and projector fixed and train only the LLM. The batch size is set to 128 with learning rate of $2\times 10^{-5}$ and warm up ratio is 0.03 with cosine learning rate scheduler.
4.2.4 GPU Burn
@@
To evaluate the external validity and real-world applicability of our experimental results, we utilized the open-source GPU Burn tool [28]. A GPU burn is designed to stress test multi-GPU setups by maximizing GPU memory utilization and compute load while validating calculation correctness. The tool forks a process for each GPU, allocates a high percentage of available memory, and performs continuous matrix multiplications using the CUBLAS library. We obtained the gpu-burn source code from the official GitHub repository at github.com/wilicc/gpu-burn and compiled it on our target Linux system following the provided instructions.
@@
4.3 Power Measurement
The GPU power draw was measured using the Weights & Biases (wandb) [29] developer interface. The DCIM, provided by NLYTE software, is used to manage, monitor and optimize the data center’s power and cooling infrastructure. Power and current data from the rack mounted PDU is pulled into the DCIM system over the network using communication protocols like SNMP and Modbus. The DCIM collects and processes this data, allowing operators to monitor, analyze, and optimize power usage at the rack level. The DCIM reports real-time data, including metrics like power draw, and alerts for potential issues related to critical infrastructure in the data center.
@@ 5 Results
Table 1: Power sampled at 1-minute intervals over training duration. Batch size in images/tokens for ResNet/Llama models.
Training Model Median Batch Duration Total Avg. GPU
Workload Task Power (kW) Size (min) Energy (kWh) Util. (%)
@@
GPU Burn (no CPU) Stress 7.97 N/A 60 8 100
GPU Burn (+CPU) Stress 8.43 N/A 77 10.5 100
6 Discussion